OpenStudy (anonymous):

Medal Given! let h(x)=6^2 and g(x)=4x+8 what is the product of h(-1) and g(-3). Answer Choices: -24 24 -120 120

OpenStudy (fibonaccichick666):

So first, what is h(any number)

OpenStudy (anonymous):

-1?

OpenStudy (fibonaccichick666):

Can you tell me how you came to that conclusion?

OpenStudy (anonymous):

its in the question?

OpenStudy (anonymous):

let h(x)=6^2 and g(x)=4x+8 what is the product of h(-1) and g(-3)

OpenStudy (anonymous):



OpenStudy (anonymous):

let h(x)=6x^2 and g(x)=4x+8 what is the product of h(-1) and g(-3)

OpenStudy (fibonaccichick666):

oh, wait no no ok so I'm just asking, does h depend on any variable?

OpenStudy (fibonaccichick666):

because as stated \[h(x)=6^2\] so is it a constant function is my question